Using a TI-84:

Jennifer Jenson learns that the Newport Temp Agency has hired only 15 women among its last 40 new employees. She also learns that the pool of applicants is very large, with an equal number of qualified men and women. Find the probability that among 40 such applicants, the number of women is 15 or fewer. Based on the result, is there strong evidence to charge that the Newport Temp Agency is discriminating against women?
Answer #1

Solution:

Given that x = 15, n = 40

p = 0.5, q = 1-p = 1- 0.5= 0.5
p-hat = 15/40 = 0.375

test statistic:
zo = (p-hat - p) /sqrt(p(1-p)/n)
zo = (0.375 - 0.5)/sqrt(0.5*(1-0.5)/40)
zo = -1.5811

zo of -1.5811 correspsonds to a p-value
P-value = 0.05695

P-value is not unusual. So there is not strong evidence to charge that the Newport temp agency is discrimating against women.